OpenStudy (anonymous):

Factor completely: x^2 − 8x + 16

OpenStudy (anonymous):

lol didnt i give you the blue prints

OpenStudy (anonymous):

Note that (a-b)^2=a^2-2ab+b^2.

OpenStudy (anonymous):

@AnimalAin thats wrong

OpenStudy (anonymous):

lol x^2 − 8x + 16 = (x-4)(x-4)=(x-4)^2

OpenStudy (anonymous):

Not the answer, it was a hint.

OpenStudy (anonymous):

this is what i got (x + 4)(x − 4)

OpenStudy (anonymous):

@surdawi is that write?

OpenStudy (anonymous):

No. surdawi solved the problem correctly above. When you have an answer in the form you do, the product will be the difference of squares, without a middle term.

OpenStudy (anonymous):

(x + 4)(x + 4) (x − 4)(x − 4) (x + 4)(x − 4) (x − 2)(x − 8)

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

You don't listen very well. The correct answer is above.

OpenStudy (snuggielad):

I see you have only been online for 4 days! Welcome to OpenStudy! Don't use this site to cheat. This is not a place for that. I also reccomend that you answer a TON of questions. People are more likely to help someone who has more answers than questions. You on the other hand have asked 183 questions and given 7 answers in only four days. Let me ask you...would you help you?

OpenStudy (anonymous):

yes i would :)

OpenStudy (snuggielad):

:| I wouldnt. I am sorry. In only four days you have managed to ask 183 questions. Hmmm...makes you kind of feel used doesnt it guys...she isnt even paying it forward.

OpenStudy (anonymous):

some people like to give just to do it because their kind. they know that some people have a hard time in math. so thank you very much

OpenStudy (anonymous):

u dnt always have to get something back

OpenStudy (anonymous):

and i would help people if i knew the answers but i dnt i suck in math!!!!! so stop juging and stop talking to me ..bye.

OpenStudy (anonymous):

if some people dnt like that i ask alot of questions dont help me plain and simple. point blan period

OpenStudy (anonymous):

im not using any body they help me threw it and show me how to do the work. only sometimes do they acually just go out and say it. soo get your facts straight

OpenStudy (anonymous):

your the selfish one not understanding that sometimes people just dnt get math at alll!!!!!!!!! thats me

OpenStudy (snuggielad):

well thats good. I am glad they dont give you answers. but you really do need to answer other peoples :) like i said pay it forward

OpenStudy (anonymous):

like i said before im not good at math and im not gonna give people answers that im not 100% sure on

OpenStudy (evonhowell):

Snuggie enough your suppose to be an ambassador and your calling someone selfish if needed Preetha will be called be the Ambassador your Suppose to be this is not ambassador like. @SnuggieLad

OpenStudy (anonymous):

CLose this question @allie_bear22

OpenStudy (anonymous):

Right now

OpenStudy (anonymous):

before things get worse

OpenStudy (anonymous):

i never knew an ambassador could be so mean i just wanted help

OpenStudy (anonymous):

now im almost afraid to post anything else

OpenStudy (anonymous):

well u have asked a lot of questions

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

srry i thin k its kinda un bliveable to

OpenStudy (anonymous):

*think it's

OpenStudy (anonymous):

Don't be afraid....just don't cheat on your hoemwork

OpenStudy (anonymous):

its not home wrok its her schoolwork

OpenStudy (yrelhan4):

Alright, the question is closed guys. Everyone get back to work. :P